OpenStudy (anonymous):

Solve by completing the square: x^2-6x+5=0

OpenStudy (anonymous):

x^2 -6x +9 = -5 +9

OpenStudy (anonymous):

What exactly does completing the square mean? What does it involve? What steps?

OpenStudy (anonymous):

(x-3)^2 = 4

OpenStudy (anonymous):

x^2+2px= x^2+2px+p^2 -p^2 = (x+p9^2 -p^2

OpenStudy (anonymous):

Alisha101, where did the nines come from?

OpenStudy (anonymous):

x^2+2px= x^2+2px+p^2 -p^2 = (x+p)^2 -p^2 Typo

OpenStudy (anonymous):

thats the completing the square formula you move the number to the other side and complete the square

OpenStudy (anonymous):

take the square root of 4 now to get -2 or 2

OpenStudy (anonymous):

solve both equations x-3 = -2 or x-3 = 2

OpenStudy (anonymous):

x = 1 or x = 5

OpenStudy (anonymous):

estudier, What... is.... THAT!?!? Haha

OpenStudy (anonymous):

That is the general procedure for completing the square without messing around with all those numbers.

OpenStudy (anonymous):

move the 5 to other side so you have -5 now you need to find the perfect square completion to x^2 -6x + _____

OpenStudy (anonymous):

if you put 9 in you can get the formula (x-3)^2 so you add 9 to both sides (what you do to one side you do to the other side

OpenStudy (anonymous):

now condense and solve

OpenStudy (anonymous):

Wouldn't you subtract 9?

OpenStudy (anonymous):

estudier, what does the p represent in your formula?

OpenStudy (anonymous):

In this case, -3

OpenStudy (anonymous):

2*-3 = -6

OpenStudy (anonymous):

Did you say there was a typo in that formula?

OpenStudy (anonymous):

This post "x^2+2px= x^2+2px+p^2 -p^2 = (x+p)^2 -p^2 Typo" corrected the one two posts earlier.

OpenStudy (anonymous):

So why is p=-3? Just so I know for different problems.

OpenStudy (anonymous):

The problem will typically be presented as x^2 + bx... p=b/2 The other thing u need to watch is if the problem is presented as ax^2 + when u need to divide through by a first.

OpenStudy (anonymous):

That makes sense! Now just one more question! So for the formula, I solve for x, but do I write it all out at once? So it would be..... x^2+2(-3)x=x^2+2(-3)+(-3)^2-(-3)^2=(x+(-3)^2-(-3)^2????? How would I solve it all at once? Or do I take the formula apart somehow?

OpenStudy (anonymous):

Also, for the option without the formula, I got x=(+-) 5

OpenStudy (anonymous):

x^2+2(-3)x=x^2+2(-3)+(-3)^2-(-3)^2=(x+(-3)^2-(-3)^2????? x^2+2(-3)x=(x+(-3)^2-(-3)^2 = (x-3)^2 - 9 After u have done a half dozen or so, u will be able to just do the thing without the middle bits on automatic...

OpenStudy (anonymous):

Also, for the option without the formula, I got x=(+-) 5 So when u add the 5 u have (x-3)^2 - 9 + 5 = (x-3)^2 -4 = 0 so so x-3 = plus or minus 2. x = -5 or 1.

OpenStudy (anonymous):

Wouldn't it be positive 5? And thanks! I forgot about the (plus minus) part!

OpenStudy (anonymous):

Yes, 5, sorry (as u may have noticed, my attention wanders with computations).

OpenStudy (anonymous):

WOW, I ACTUALLY GOT IT! Thank you! :)

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

:-D